Hello, how many terms of the sequence -19,-15,-11,... must be taken to give a sum of 98? Thanks :)
Answer by stanbon(75887)

how many terms of the sequence -19,-15,-11,... must be taken to give a sum of 98?
---
a = -19
d = 4
--------
S(n) = (n/2)(a + L) where L is the nth term:
L = a + (n-1)d
================
Solve:
98 = (n/2)(-19 + -19 +(n-1)4)
196 = n(-38+4n-4)
196 = n(4n-42)
98 =n(2n-21)
2n^2-21n-98 = 0
(n-14)(2n+7) = 0
---
Positive solution:
n = 14
Ans: 14 terms
=================
Cheers,
Stan H.